# MCGrad: Multicalibration at Web Scale

### v2

@@ 4.2. Experimental Results
Table 1 reports the average performance over the datasets. MCGrad achieves by far the best average log loss improvement ($10.4\%$) and PRAUC improvement ($8.1\%$) over the base predictor. Besides these average improvements over datasets MCGrad never harms performance which helps build the necessary trust among ML engineers to delpoy it in production environments. Other methods don’t have this guarantee. For instance while, DFMC improves model performance metrics by a lower $1.8\%$ and $0.76\%$, it harms PRAUC and log loss on two datasets. Isotonic and HKRR even decrease the PRAUC on average by, respectively, $1\%$ and $5\%$ with a negligible change in log loss. For each metric, we rank all compared methods from $1$ (best) to $5$ (worst) and report the average ranking position in Table 1. A lower average rank indicates how often the method is preferred, as opposed to the average performance, which reflects the magnitude. MCGrad consistently ranks as #1\#1 baseline on log loss and PRAUC, while only ranking third for MCE on the MEPS dataset. Finally, MCGrad ranks second best method in terms of global calibration (ECCE), losing against Isotonic Regression, which is expected because it is trained to directly optimize the global calibration of the base predictor.
@@
Because MCGrad recurrently improves its own predictions by using $T$ rounds, one relevant question is whether $T>1$ is really necessary in practice. For this goal, we compare MCGrad with its variant Only One Round that forces the number of rounds T=1T=1.333Early stopping can possibly return $T=0$. Figure 2 (purple line) shows the relative improvement of Only One Round for the four considered metrics and all benchmark datasets. In the majority of cases, it is evident that Only One Round deteriorates MCGrad’s performance: on the MCE and ECCE axes, it more than doubles its value (relative improvement <100<-100) for, respectively, $6$ and $5$ datasets, while, on the log loss and PRAUC axes, it worsens the performance by approximately $2\%$ or more on, respectively, $8$ and $6$ out of $11$ datasets. Interestingly, the largest deterioration of performance corresponds to MCGrad running for the highest number of rounds: the optimal $T$ for ACSEmploy, ACSIncome, ACSPublic, and ACSMobility is, respectively, $20$, $35$, $27$ and $22$.
@@
Rescaling the logits of all predictions by learning a constant speeds up MCGrad training by reducing rounds, especially on very large datasets. As a downside, it might affect MCGrad’s performance. Figure 2 (light blue bar) shows the relative deterioration after removing rescaling. Effects are mild and balanced: removal worsens MCE, log loss, and PRAUC on $6$, 66, and $7$ datasets respectively, with average drops of $9\%$, $0.5\%$, and $0.03\%$.
@@ 5. Results in Production Deployment
We describe the results obtained by using MCGrad in a real-world production system. MCGrad has been deployed and used in production at Meta. Specifically, we have integrated MCGrad into two ML training platforms at Meta: the publicly released Looper (Markov et al., 2022), and an internal platform that we refer to as MLplatform#2. The evaluation consists of a total of $147$ binary classification models that serve live production traffic: $27$ models on Looper and $120$ models on the other platform. On both ML training platforms, we run MCGrad with default hyperparameters (see Appendix C).
@@
For 27 binary classification models that serve active production traffic on Looper, we created a variant of the model that applies MCGrad as calibration post-processing. We started online A/B tests to compare these 27 models against the active production models, all of which are GBDTs. Those production models all had previously already applied Platt scaling (Platt et al., 1999).
We found that on $24$ out of $27$ models, the variant with MCGrad statistically significantly outperformed the same model with Platt scaling (Platt et al., 1999), resulting in the promotion of the MCGrad-calibrated model variant to become the primary production model.
@@
This ML training platform is widely used at Meta. ML model training scripts on this platform define the training set, a validation set, the test set, the ML model architecture, and the set of model evaluation metrics. The ML platform allows running an evaluation flow that trains the model on the training set and evaluates the model on the test set using the specified set of evaluation metrics, or an publish flow that prepares an API endpoint that can generate predictions for live production traffic. We integrated MCGrad into the evaluation flow so that we obtain an additional set of evaluation results for a variant of the model that is calibrated using MCGrad.
For every ML model evaluation on this ML platform, we collected parallel evaluation results with and without MCGrad. We conducted data collection in two separate periods. Table 3 summarizes the results. In the first data collection period, we obtained results on $120$ binary classification production models. We found that applying MCGrad post-processing improved logloss for $88.7\%$ of the models, compared to the production models without it. PRAUC improved for $76.7\%$ of the models, AUROC for $80.2\%$, and Expected Calibration Error (ECE) (Naeini et al., 2015) for $86\%$. Note that on most models where no improvement was found, there was no metric degradation either, due to the early stopping. In Table 3 we can see that that the worst 10th-percentile metric effects of MCGrad still result in degradations. We manually investigated a sample of those models, and found that these were mostly explained by errors by ML practitioners in the train and test set specification. E.g., in some cases the train and test set were clearly from a different population, or sample weights were defined in the train set but not in the test set.
@@ 6. Learnings from Production Deployment
Our results demonstrate that multicalibration post-processing can improve real-world models’ multicalibration as well as performance metrics that tend to correlate with business outcomes, such as PRAUC. Furthermore, baselines such as Isotonic Regression do not achieve similar results. This learning contradicts some observations by Hansen et al. (2024): (1) Isotonic Regression is often competitive compared to algorithms designed specifically for multicalibration, and that (2) practitioners may face a trade-off between multicalibration and predictive performance. MCGrad achieves better results than baselines and rarely degrades performance.
@@
Many multicalibration algorithms (Hebert-Johnson et al., 2018; Jin et al., 2025) require manual specification of the protected groups. In our experience deploying multicalibration in production, we found that many ML engineers struggle to manually define relevant protected groups, and lack guidelines and frameworks for defining them. This experience is in line with findings from a series of interviews on AI fairness adoption in industry (Holstein et al., 2019), where it turns out that practitioners often lack knowledge on different types of biases and cannot identify the relevant biases to correct for. Hence, our experience is that in order to succeed in obtaining industry adoption, a multicalibration method needs to relieve users of the burden of manually defining protected groups, like MCGrad does.
Learning 3. Practitioners consider adopting multicalibration methods only when they work out-of-the-box, reliably, and with no risk of harming model performance
Many multicalibration algorithms have hyperparameters. We found that industry practitioners are open to applying multicalibration on their production ML models, but only if the required effort is limited. Hyperparameter tuning, for example, is often seen as too complex or time-consuming. Furthermore, they perceive it as a reliability risk: if a method requires hyperparameter tuning to work well, then they wonder if the hyperparameter configuration will still work well in subsequent model training runs.
@@
We defined a set of hyperparameter default values for MCGrad that achieves (1) no degradation in PRAUC and log loss, and (2) substantial reduction in MCE. Appendix C details the procedure used to obtain these hyperparameters. Results in Sections 4 and 5 confirm that those default values work sufficiently well to consistently produce log loss and PRAUC improvements over models calibrated with Platt scaling (on Looper) or uncalibrated models (on MLplatform#2), and consistently yield log loss, PRAUC, and MCE improvements on the public datasets.
@@ 7. Related Work
Multicalibration. Multicalibration as a learning objective was first introduced by (Hebert-Johnson et al., 2018), requiring a predictor to remain calibrated even when conditioned on membership in any of a potentially large collection of subgroups of the domain and the predicted value.
Considerable work has been devoted to developing the theoretical foundations of multicalibration, focusing on computational complexity and elucidating its relation to other learning concepts (Detommaso et al., 2024; Gopalan et al., 2023; Deng et al., 2023; Błasiok et al., 2024).
Low-degree multicalibration (Gopalan et al., 2022b) generalizes conditioning on the prediction by using a weight function of the predicted value belonging to a class of weight functions. It was shown that lower computational complexity can be achieved by restricting the class of weight functions to low-degree polynomial functions.
Swap multicalibration (Gopalan et al., 2023) is a stronger concept that implies multicalibration. This stronger condition has been shown to be satisfied by some existing multicalibration algorithms, such as the algorithm of (Hebert-Johnson et al., 2018). Swap multicalibration has been shown to be equivalent to notions of swap agnostic learning and swap omniprediction.
Omnipredictors, a notion originally introduced in (Gopalan et al., 2022a), strengthen the classical notion of agnostic learning by requiring a predictor to provide a minimization guarantee for each loss function in a class of loss functions. This concept was extended in (Gopalan et al., 2023) to the notion of swap omnipredictors. (Gopalan et al., 2023) has shown equivalence between swap multicalibrated predictors and swap omnipredictors for the class of convex functions satisfying certain conditions.
In (Haghtalab et al., 2023b), a connection between multicalibration and multi-objective learning is noted and exploited to formulate the learning as a game dynamics to achieve state-of-the-art guarantees for a diverse set of multicalibration learning problems. Recent work (Okoroafor et al., 2025) studied omniprediction algorithms and their theoretical guarantees.
@@
Algorithms for multicalibration. The early proposals of multicalibration algorithms, such as HKRR (Hebert-Johnson et al., 2018) (e.g., implemented in MCBoost (Pfisterer et al., 2021) or by Hansen et al. (2024)) and LSBoost (Globus-Harris et al., 2023), rely on discretizing the output space and calibrating the prediction for each discretization level until convergence. A discretization-free multicalibration algorithm was proposed in (Jin et al., 2025), requiring group functions as input and using a gradient boosting machine to fit an ensemble of depth-two decision trees to the residuals of an uncalibrated predictor using both the original features and the predictions of the uncalibrated predictor. For guaranteeing multicalibration within an additive error, this algorithm requires a loss saturation condition to hold, i.e., that the squared-error loss can be further reduced only by a small amount. Our algorithm uses gradient boosting in a similar manner, but does not require pre-specified groups, and leverages a multi-round recursive gradient boosting and allows for arbitrary user-defined decision trees to be used in each round.
8. Conclusion
@@
We tackle the problem of designing a multicalibration algorithm without having access to pre-specified protected sets, that is fast and safe to deploy in production. These three aspects are relevant for practitioners as they are often, (1) unable to define good groups for their task, (2) forced to limit the overhead in terms of memory and speed, and (3) required to guarantee that the algorithm does not harm the base predictor. We proposed MCGrad, a novel algorithm that achieves multicalibration by optimally finding the regions of the feature space where the miscalibration is large, and fixing the predictions. Because this might create new regions with evidence of miscalibration, MCGrad employs multiple rounds, where in each round it corrects its own output. In addition, MCGrad’s implementation is fast, lightweight, and safe to deploy. We showed that MCGrad outperforms existing baselines on benchmark datasets and provided results obtained in a large-scale industry production deployment at Meta consisting of hundreds of production models. Finally, we included a set of practical takeaways from our experience applying multicalibration in industry.
Appendix A Theoretical Results
A.1. On the Multicalibration Error
@@
Different definitions of approximate multicalibration have been considered, referred to as $\alpha$-multicalibration. One common choice is using a uniform bound on the absolute deviation, $\Delta_{h,g}(f)\leq\alpha$ where the precise choice of $h,g$ varies (Błasiok et al., 2024; Deng et al., 2023; Gopalan et al., 2022b; Haghtalab et al., 2023a). Another common choice is a uniform bound on the conditional deviation, $\delta_{h,g}(f):=|\mathbb{E}\left[h(X)g(f(X))(Y-f(X))\mid h(X)g(f(X))=1\right]|\leq\alpha$, (Hebert-Johnson et al., 2018), (Shabat et al., 2020). These definitions were improved on by (Haghtalab et al., 2023a) which showed that a tighter bound can be achieved with no additional requirements (”for free”) by defining $\alpha$-multicalibration as $\Delta_{h,g}(f)\leq\alpha\sqrt{\mathbb{E}\left[h(X)\right]}$.
@@
and the scale parameter chosen as
(4) $\tau_{h}(f)=\sqrt{\mathbb{E}_{D}\left[h(X)f(X)(1-f(X))\right]},$
matches with the MCE metric introduced in Guy et al. (2025), defined in Equation (2).
@@
Given a set $\mathcal{H}$ of group membership functions, a probabilistic predictor $f$ is $\alpha$-multicalibrated with respect to $\mathcal{H}$, with the scale parameter $\tau_{h}(f)$ as in (4), if and only if $\text{MCE}(f)\leq\alpha\sqrt{n}$.
@@
First, note that
$\Delta_{h,g}(f)=\frac{1}{n}\big|\sum_{k=1}^{n}h(x_{k})g(f(x_{k}))(y_{k}-f(x_{k}))\big|.$
Combining with the definition of $\text{ECCE}_{h}(f)$, we get that
@@
Plugging in the definition of $\text{MCE}(f)$, given in Equation (2), we have
$\frac{1}{\sqrt{n}}\text{MCE}(f)$ $=$ $\frac{1}{\sqrt{n}}\max_{h\in\mathcal{H}}\frac{\text{ECCE}_{h}(f)}{\sigma_{h}(f)}$
$=$ $\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\sqrt{n}\Delta_{h,g}(f)}{n_{h}\sigma_{h}(f)}$
$=$ $\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\Delta_{h,g}(f)}{\tau_{h}(f)}$
@@
A.2. Convergence Analysis
In this section, we analyze the convergence of Algorithm 1 under certain conditions for the GBM used in each round of the algorithm. Specifically, we show two results. First, we show conditions under which the loss is non-decreasing with the number of rounds, and is strictly decreasing unless the predictor has converged. Second, we show conditions under which the $\alpha$ multicalibration error after $T$ rounds of the algorithm is bounded in terms of the gap between predictors $F_{T+1}$ and $F_{T}$, achieving 0-multicalibration when the predictor has converged.
The results in this section account for the fact that, in each round $t$, the GBM iteratively adds regression trees to $F_{t}$ to form $F_{t+1}$. This is different from the arguments made in Section 3.1 where we assume that an optimal linear combination of weak learners, minimizing the loss function, is added in each round of the algorithm.
We consider a GBM that in round $t$ recursively combines $M_{t}$ regression trees as follows:
$F_{t}^{m+1}(x)=F_{t}^{m}(x)+\rho_{m}h_{j_{m}}(x,f_{t}(x))$
where $j_{m}$ is the index of the regression tree added in the $m$-th iteration, $\rho_{m}$ is the step size, and $m=0,1,\ldots,M_{t}-1$.
Let

### v3

@@ 4.2. Experimental Results
Figure 3. Computational time of multicalibration algorithms with varying numbers of groups. (Left) Fit time and (Right) predict time in seconds as a function of the number of groups on a log-log scale. Runtime for DFMC and HKRR increases with the number of groups, while MCGrad is constant.
@@
Table 2 reports the average performance over the datasets. MCGrad achieves by far the best average log loss improvement ($10.4\%$) and PRAUC improvement ($8.1\%$) over the base predictor. Besides these average improvements over datasets, MCGrad never harms performance, which helps build the necessary trust among ML engineers to deploy it in production environments. Other methods don’t have this guarantee. For instance, while DFMC improves model performance metrics by a lower $1.8\%$ and $0.76\%$, it harms PRAUC and log loss on two datasets. Isotonic and HKRR even decrease the PRAUC on average by, respectively, $1\%$ and $5\%$ with a negligible change in log loss. For each metric, we rank all compared methods from $1$ (best) to $5$ (worst) and report the average ranking position in Table 2. A lower average rank indicates how often the method is preferred, as opposed to the average performance, which reflects the magnitude. MCGrad consistently ranks as #1\#1 baseline on log loss and PRAUC, while only ranking third for MCE on the MEPS dataset. Finally, MCGrad ranks second best method in terms of global calibration (ECCE), losing against Isotonic Regression, which is expected because it is trained to directly optimize the global calibration of the base predictor.
@@
Because MCGrad recurrently improves its own predictions by using $T$ rounds, one relevant question is whether $T>1$ is really necessary in practice. For this goal, we compare MCGrad with its variant Only One Round that forces the number of rounds T=1T=1.555Early stopping can possibly return $T=0$. Figure 2 (purple line) shows the relative improvement of Only One Round for the four considered metrics and all benchmark datasets. In the majority of cases, it is evident that Only One Round deteriorates MCGrad’s performance: on the MCE and ECCE axes, it more than doubles its value (relative improvement <100<-100) for, respectively, $6$ and $5$ datasets, while, on the log loss and PRAUC axes, it worsens the performance by approximately $2\%$ or more on, respectively, $8$ and $6$ out of $11$ datasets. Interestingly, the largest deterioration of performance corresponds to MCGrad running for the highest number of rounds: the optimal $T$ for ACSEmploy, ACSIncome, ACSPublic, and ACSMobility is, respectively, $20$, $35$, $27$ and $22$.
@@
Rescaling the logits of all predictions by learning a constant speeds up MCGrad training by reducing rounds, especially on very large datasets. As a downside, it might affect MCGrad’s performance. Figure 2 (light blue) shows the relative deterioration after removing rescaling. Effects are mild and balanced: removal worsens MCE, log loss, and PRAUC on $6$, 66, and $7$ datasets respectively, with average drops of $9\%$, $5\%, 0.0$$03\%.$
@@
Q5. Computational time of MCGrad vs multicalibration baselines as the number of groups scales.
Scalability can be a relevant bottleneck for multicalibration deployment in production environments. To evaluate this, we measure the training (fit) and inference (predict) times of MCGrad, HKRR, and DFMC as the number of protected groups increases from $10$ to $1000$, using $\approx 1M$ training and $\approx 300k$ test samples from the ACSEmploy dataset.
Figure 3 shows the methods’ computational time per sample (in microseconds) as a function of the number of groups. Notably, the gap between MCGrad and each baseline becomes larger as group count increases: at $1000$ groups, MCGrad maintains nearly constant inference time ($\sim 20\mu s$) while HKRR and DFMC require, respectively, $5019$ and $1292$ $\mu s$, meaning MCGrad achieves less than $2\%$ of their latency overhead. This gap becomes even larger in production settings where models with $100$s of features can generate tens of thousands of potential protected groups.
More importantly, these seemingly small latency differences have strong business impact. Production models typically serve online inference where even millisecond-scale increases in response time negatively affect the system performance and the user experience. Existing research claims that an increase of latency directly correlates with reduced user engagement and online business metrics (Kohavi et al., 2014). MCGrad’s sub-$20\mu s$ inference overhead allows deployment in production, while baseline methods requiring few milliseconds per prediction can be impractical for production use.
@@ 5. Results in Production Deployment
We describe the results obtained by using MCGrad in a real-world production system. MCGrad has been deployed and used in production at Meta. Specifically, we have integrated MCGrad into two ML training platforms at Meta: the publicly released Looper (Markov et al., 2022), and an internal platform that we refer to as MLplatform#2. The evaluation consists of a total of $147$ binary classification models that serve live production traffic: $27$ models on Looper and $120$ models on the other platform. On both ML training platforms, we run MCGrad with default hyperparameters.
@@
For 27 binary classification models that serve active production traffic on Looper, we created a variant of the model that applies MCGrad as calibration post-processing. We started online A/B tests to compare these 27 models against the active production models, all of which are GBDTs. Those production models all had previously already applied Platt scaling (Platt, 1999).
We found that on $24$ out of $27$ models, the variant with MCGrad statistically significantly outperformed the same model with Platt scaling (Platt, 1999), resulting in the promotion of the MCGrad-calibrated model variant to become the primary production model.
@@
This ML training platform is widely used at Meta. ML model training scripts on this platform define the training set, a validation set, the test set, the ML model architecture, and the set of model evaluation metrics. The ML platform allows running an evaluation flow that trains the model on the training set and evaluates the model on the test set using the specified set of evaluation metrics, or a publish flow that prepares an API endpoint that can generate predictions for live production traffic. We integrated MCGrad into the evaluation flow so that we obtain an additional set of evaluation results for MCGrad.
For every ML model evaluation on this ML platform, we collected parallel evaluation results with and without MCGrad. We conducted data collection in two separate periods. Table 4 summarizes the results. In the first data collection period, we obtained results on $120$ binary classification production models. We found that applying MCGrad post-processing improved log loss for $88.7\%$ of the models, compared to the production models without it. PRAUC improved for $76.7\%$ of the models, AUROC for $80.2\%$, and Expected Calibration Error (ECE) (Naeini et al., 2015) for $86\%$. Note that on most models where no improvement was found, there was no metric degradation either, due to the early stopping. In Table 4 we can see that the worst 10th-percentile metric effects of MCGrad still result in degradations. We manually investigated a sample of those models, and found that these were mostly explained by errors by ML practitioners in the train and test set specification. E.g., in some cases the train and test set were clearly from a different population, or sample weights were defined in the train set but not in the test set.
@@ 6. Learnings from Production Deployment
Our results demonstrate that MCGrad can improve real-world models’ multicalibration as well as performance metrics that tend to correlate with business outcomes, such as PRAUC. Furthermore, baselines such as Isotonic Regression do not achieve this. This contradicts observations by Hansen et al. (2024) that (1) Isotonic Regression is often competitive compared to algorithms designed specifically for multicalibration, and (2) practitioners may face a trade-off between multicalibration and predictive performance.
@@
Many multicalibration algorithms (Hebert-Johnson et al., 2018; Jin et al., 2025) require manual specification of the protected groups. In our experience deploying multicalibration in production, we found that many ML engineers struggle to manually define relevant protected groups, and lack guidelines and frameworks for defining them. This experience is in line with findings from a series of interviews on AI fairness adoption in industry (Holstein et al., 2019), where it turns out that practitioners often lack knowledge on different types of biases and cannot identify the relevant biases to correct for. Hence, our experience is that in order to succeed in obtaining industry adoption, a method needs to relieve users of the burden of manually defining protected groups, like MCGrad does.
Learning 3. Practitioners consider adopting multicalibration methods only when they work out-of-the-box, reliably, and with no risk of harming model performance.
Many multicalibration algorithms have hyperparameters. We found industry practitioners to be open to applying multicalibration to production ML models, but only if it takes limited effort. Hyperparameter tuning is often seen as complex or time-consuming. They also perceive it as a reliability risk: if a method requires hyperparameter tuning to work well, then they wonder if the hyperparameter configuration will still work well in subsequent training runs.
@@
We defined a set of hyperparameter default values for MCGrad that achieves (1) no degradation in PRAUC and log loss, and (2) substantial reduction in MCE. By pre-computing defaults through meta-analysis across $35$ datasets, we enable MCGrad to deliver consistent improvements without requiring domain expertise or extensive experimentation. Results in Sections 4 and 5 confirm that those default values consistently produce log loss and PRAUC improvements over models calibrated with Platt scaling (on Looper) or uncalibrated models (on MLplatform#2), and consistently yield log loss, PRAUC, and MCE improvements on the public datasets.
@@ 7. Related Work
Multicalibration was introduced by Hebert-Johnson et al. (Hebert-Johnson et al., 2018) as a learning objective requiring predictors to be calibrated across a large collection of subgroups. Subsequent work has deepened its theoretical foundations, exploring computational complexity and connections to other learning paradigms (Detommaso et al., 2024; Gopalan et al., 2023; Deng et al., 2023; Błasiok et al., 2024).
Extensions such as low-degree multicalibration (Gopalan et al., 2022b) generalize subgroup definitions using weight functions, showing that restricting to low-degree polynomials can reduce computational complexity. Swap multicalibration (Gopalan et al., 2023) strengthens the original concept and is satisfied by some existing algorithms; it is also equivalent to swap omniprediction for certain loss function classes. The notion of omnipredictors (Gopalan et al., 2022a) further generalizes agnostic learning by requiring minimization guarantees for a class of loss functions (Okoroafor et al., 2025). Connections to multi-objective learning have also been explored, framing multicalibration as a game dynamics problem (Haghtalab et al., 2023).
@@
Algorithms for multicalibration. The early proposals of multicalibration algorithms, such as HKRR (Hebert-Johnson et al., 2018) (e.g., implemented in MCBoost (Pfisterer et al., 2021) or by Hansen et al. (2024)) and LSBoost (Globus-Harris et al., 2023), rely on discretizing the output space and calibrating the prediction for each discretization level until convergence. A discretization-free multicalibration algorithm was proposed in (Jin et al., 2025), requiring group functions as input and using a gradient boosting machine to fit an ensemble of depth-two decision trees to the residuals of an uncalibrated predictor using both the original features and the predictions of the uncalibrated predictor. For guaranteeing multicalibration within an additive error, this algorithm requires a loss saturation condition to hold, i.e., that the squared-error loss can be further reduced only by a small amount. Our algorithm uses gradient boosting similarly but does not require pre-specified groups. Instead, it leverages multi-round recursive gradient boosting and supports arbitrary user-defined decision trees in each round.
8. Conclusion & Limitations
@@
We tackle the problem of designing a multicalibration algorithm without having access to pre-specified protected sets, that is fast and safe to deploy in production. These three aspects are relevant for practitioners as they are often (1) unable to define groups for their task, (2) forced to limit the overhead in terms of memory and speed, and (3) required to guarantee that the algorithm does not harm the base predictor. We proposed MCGrad, a novel algorithm that achieves multicalibration by finding and fixing regions of the feature space where the miscalibration is large. Because this might create new regions with evidence of miscalibration, MCGrad employs multiple rounds, where in each round it corrects its own output. In addition, MCGrad’s implementation is fast, lightweight, and safe to deploy. We showed that MCGrad outperforms existing baselines on benchmark datasets and provided results obtained in a large-scale industry production deployment at Meta consisting of hundreds of production models. Finally, we included practical takeaways from our experience applying multicalibration in industry.
Limitations.
Although MCGrad shows improvements, several limitations remain. First, guardrails against harming the base predictor may not be sufficient under concept drift. Specifically, MCGrad can potentially worsen performance on post-drift data distributions, as multicalibration is only robust to certain types of distribution shift (Roth, 2022; Kim et al., 2022). Second, MCGrad is currently limited to binary classification settings. Although multiclass extensions exist, they often suffer from reduced sample efficiency and increased complexity, creating challenges for real-world deployment (Zhao et al., 2021). Third, while MCGrad can be applied to other data modalities such as text and images (using embeddings), these applications have not been explored in this paper. Future research may study MCGrad’s (1) robustness to broader forms of drift, (2) extension to multiclass/regression problems, and (3) effectiveness on other data modalities.
Acknowledgments
We gratefully acknowledge Nicolas Stier, Udi Weinsberg, Ido Guy, Thomas Leeper, and Mark Tygert for their valuable discussions and support, which contributed to the success of this work.
Appendix A Bridging Multicalibration Error and $\alpha$-MC
@@
Several definitions of approximate multicalibration, known as $\alpha$-multicalibration, exist in the literature. A common approach uses a uniform bound on absolute deviation $\Delta_{h,g}(f)\leq\alpha$ where the precise choice of $h$ and $g$ varies (Błasiok et al., 2024; Deng et al., 2023; Gopalan et al., 2022b; Haghtalab et al., 2023). Another common choice is a uniform bound on the conditional deviation (Hebert-Johnson et al., 2018; Shabat et al., 2020), i.e.
$\delta_{h,g}(f):=|\mathbb{E}\left[h(X)g(f(X))(Y-f(X))\mid h(X)g(f(X))=1\right]|\leq\alpha.$ Recent work (Haghtalab et al., 2023) improved these by showing a tighter bound $\Delta_{h,g}(f)\leq\alpha\sqrt{\mathbb{E}\left[h(X)\right]}$ can be achieved with no extra requirements.
@@
with scale parameter chosen $\tau_{h}(f)=\sqrt{\mathbb{E}_{D}\left[h(X)f(X)(1-f(X))\right]}$, matches with the MCE metric defined in Eq. (2).
@@
Given a set $\mathcal{H}$ of group membership functions, a probabilistic predictor $f$ is $\alpha$-multicalibrated with respect to $\mathcal{H}$, with the scale parameter $\tau_{h}(f)$, if and only if $\text{MCE}(f)\leq\alpha\sqrt{n}$.
@@
Note that $\Delta_{h,g}(f)=\frac{1}{n}\big|\sum_{k=1}^{n}h(x_{k})g(f(x_{k}))(y_{k}-f(x_{k}))\big|.$ Combining with the definition of $\text{ECCE}_{h}(f)$, we get that
@@
Plugging in the definition of $\text{MCE}(f)$, as in Eq. (2), we have
$\frac{1}{\sqrt{n}}\text{MCE}(f)\!=\!\frac{1}{\sqrt{n}}\max_{h\in\mathcal{H}}\!\frac{\text{ECCE}_{h}(f)}{\sigma_{h}(f)}\!=\!\!\max_{h\in\mathcal{H},g\in\mathcal{G}}\!\!\frac{\sqrt{n}\Delta_{h,g}(f)}{n_{h}\sigma_{h}(f)}\!=\!\max_{h\in\mathcal{H},g\in\mathcal{G}}\!\frac{\Delta_{h,g}(f)}{\tau_{h}(f)}.$
@@
Appendix B Convergence Analysis
We analyze the convergence of Algorithm 1 under specific conditions for the GBM used in each round. We show that (1) the loss does not increase with more rounds and decreases unless the predictor has converged, and (2) the $\alpha$-multicalibration error after $T$ rounds is bounded by the difference between predictors $F_{T+1}$ and $F_{T}$ and that 0-multicalibration is achieved at convergence. Unlike Section 3.1, which assumes adding an optimal linear combination of weak learners each round, here the GBM incrementally adds regression trees to $F_{t}$ to form $F_{t+1}$.
We consider a GBM that, in round $t$, iteratively combines $M_{t}$ regression trees as $F_{t}^{m+1}(x)=F_{t}^{m}(x)+\rho_{m}h_{j_{m}}(x,f_{t}(x))$ where $j_{m}$ is the index of the regression tree added in the $m$-th iteration, $\rho_{m}$ is the step size, and $m=0,1,\ldots,M_{t}-1$. Let